OpenStudy (anonymous):

Please help. Kristen can run 15km h^-1 and walk at 5km h^h-1. She completed a 42 km marathon in 4 hours. What distance did Kristen run during the marathon?

OpenStudy (anonymous):

15x + 5y = 42 x + y = 4 (times this by five so you can get rid of one value) 15x + 5y = 42 - 5x + 5y = 20 = 10x = 22 x = 2.2 (sub back into original) y = 1.8

OpenStudy (whpalmer4):

Of course, this all assumes she didn't spend any of that 4 hours at any other speed, or stopped, which isn't very realistic :-) I agree with jordirh's result, however.